Find the p-values for the following critical values: (Assume two sided hypothesis) 2.03 1.50 1.40 2.26

  1. Find the p-values for the following critical values: (Assume two sided hypothesis)
    1. 2.03
    2. 1.50
    3. 1.40
    4. 2.26
Answer #1

a)

p-value = 2 * P(Z > z)

= 2 * P(Z > 2.03)

= 2 * 0.0212

= 0.0424

b)

p-value = 2 * P(Z > z)

= 2 * P(Z > 1.50)

= 2 * 0.0668

= 0.1336

c)

p-value = 2 * P(Z > z)

= 2 * P(Z > 1.40)

= 2 * 0.0808

= 0.1616

d)

p-value = 2 * P(Z > z)

= 2 * P(Z > 2.26)

= 2 * 0.0119

= 0.0238